OpenStudy (anonymous):

How do I solve using the substitution method? y = 2x - 5 3y - x = 5

OpenStudy (anonymous):

Since y = 2x - 5, you can put substitute that into the second equation: 3y - x = 5 3(2x-5) - x = 5 just solve for x from here...

OpenStudy (anonymous):

just sub 2x - 5 in for y in the second equation, then solve for x.

OpenStudy (anonymous):

3(3x-5) - x = 5 9x - 15 - x = 5 8x - 15 = 5 8x = 20 /8 /8 x = 2.5

OpenStudy (anonymous):

I made a mistake

OpenStudy (anonymous):

no.....2x - 5 not 3x - 5.....

OpenStudy (anonymous):

3(2x-5) - x = 5 6x - 15 - x = 5 5x - 15 = 5 5x = 20 x = 4 ?

OpenStudy (anonymous):

you got it James....now you can sub 4 in for x in either of the original equations and solve for y.

OpenStudy (anonymous):

omg. Thanks a lot!

OpenStudy (anonymous):

your welcome....do you know how to check your answer to see if it is correct ?

OpenStudy (anonymous):

y = 2 * 4 - 5 y = 3

OpenStudy (anonymous):

(4,3)

OpenStudy (anonymous):

Let me check because I haven't even done the problem yet.... 3(2x - 5) - x = 5 6x - 15 - x = 5 5x = 20 x = 4 y = 2x - 5 y = 2(4) - 5 y = 8 - 5 y = 3 check... 3y - x = 5 3(3) - 4 = 5 9 - 4 = 5 5 = 5 (correct) or/and check with the other equation... y = 2x - 5 3 = 2(4) - 5 3 = 8 - 5 3 = 3 (correct) ANSWER : (4,3) You got it James.....way to go :)

OpenStudy (anonymous):

seriously thanks for the help!

OpenStudy (anonymous):

not a problem...thats what I am here for :)